OpenStudy (anonymous):

18. Use the discriminant to find the number of real solutions of the equation 3x2 – 5x + 4 = 0. Show your work.

OpenStudy (keygrover):

3x2= 6 6-5x+4=0 6+4=10 -5x+10=0 0-10=-10 -5x=-10 X=2

OpenStudy (anonymous):

no is 3^2

OpenStudy (keygrover):

Oh ok

OpenStudy (anonymous):

@keygrover